OpenStudy (anonymous):

What is the monthly payment of a car for $6,500 for 4 years, at the rate of 18%. I need the equation too!

OpenStudy (wolf1728):

You can find the formula here: http://www.1728.org/loanform.htm

OpenStudy (brunsonni):

I don't use equations to solve percentages, but I got the monthly cost of $1,170, and the 4 year cost of $56,160

OpenStudy (wolf1728):

You'll need the equation. Monthly Payment = 190.94 Total paid = 9,165.12 I can use the formula and show you.

OpenStudy (wolf1728):

Where did you get your formula?

OpenStudy (wolf1728):

Monthly Payment = [rate + ((rate/((1 + rate)^months)-1)] * principal

OpenStudy (wolf1728):

rate = rate/1200 18 / 1200 = .015 Monthly Payment = [.015 + (((.015/((1.015)^48-1)] * 6,500 Monthly Payment = .015 + [.015 / 2.0434782893 -1] * 6,500 Monthly Payment = (.015 + 0.0143749996) * 6500 Monthly Payment = 0.0293749996 * 6500 Monthly Payment = 190.9374974542 which rounds to 190.94

OpenStudy (wolf1728):

Hello again Petepw, are you going to post your thoughts on this?

OpenStudy (anonymous):

is 18 the annual rate, and is it compounded?